Balls numbered from 1 to 52 are placed in a jar and mixed. If one is drawn at random, find the probability of each result.
A. The number is less than 15.
B. The number is greater than or equal to 20.
C. The number is a 10 or a 20.
D. The number is less than 11 or greater than 48.

Answer by venugopalramana(3286)

Balls numbered from 1 to 52 are placed in a jar and mixed. If one is drawn at random, find the probability of each result.
HENCE TOTAL POSSIBILITIES = 52
A. The number is less than 15.....THERE ARE 14 SUCH BALLS ..HENCE P=14/52=7/26
B. The number is greater than or equal to 20. ...THERE ARE 33 SUCH BALLS...HENCE P=33/52
C. The number is a 10 or a 20.....THERE ARE 2 SUCH BALLS....HENCE P=2/52=1/26
D. The number is less than 11 or greater than 48...THERE ARE 10+4=14 SUCH BALLS...HENCE P=14/52=7/26